1.Define potentials. Write Maxwell’s equations in term of potentials. Discuss Green function method of solving wave equations.


2.Derive and discuss electromagnetic field of a point charge in arbitrary motion. Derive and discuss total power radiated and its angular distribution by an an accelerated charge.


3.Define vectors and tensors in 4-dimentional Minkoweskiin space-time. Define 4-current density, electromagnetic field tensor and its dual and derive Maxwell’s eq 1 uations in 4-dimensional covariant form.


4.            Write down Maxwell’s equations. Show that the equation of continuity follows from Maxwell’s equations. Derive and discuss Poynting theorem and conservation law of energy and momentum for a system of charged particles and electromagnetic field

Grade:

1 Answers

Aman Bansal
11 years ago

Dear Richa,

Maxwells equations in terms of vector potential

\nabla\cdot\textbf{A} + \frac{1}{c^2} \frac{\partial \phi}{\partial t} = 0.

Using the Lorenz gauge, Maxwells equations can be written compactly in terms of the magnetic vector potential A and the electric scalar potential Φ.

\nabla^2\phi - \frac{1}{c^2}\frac{\partial^2 \phi}{\partial t^2} = - \rho/ \epsilon_0
\nabla^2\textbf{A} - \frac{1}{c^2}\frac{\partial^2 \textbf{A}}{\partial t^2} = - \mu_0 \textbf{J}

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert